# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v1

@@ 3.2.6. Distribution of Sample Sizes
C-19 (zhou2023deep, ) 2023 PD, ALS, HD Sensor Force sensitive resistor LSTM, GRU, ResNet, FCN, etc 64 75.4-83.3 0.56
Table 5. Existing studies of AI-assisted diagnosis of AD (A-1 to A-5), ALS (S-1 to S-3), HD (H-1 to H-7) or MS (M-1 to M-6) based on gait.
@@
A-1 (costa2016application, ) 2016 AD Sensor IMUs SVM, MLP, RBNs, DBSs 72 91.0-96.6 0.56
A-2 (you2020alzheimer, ) 2020 AD Sensor+Vision Camera+EEG SVM, RF, CNN, ST-GCN, etc 87 91.1-98.6 0.89
A-3 (ghoraani2021detection, ) 2020 AD Sensor Pressure sentsitive mat SVM 78 78.0 0.44
A-4 (seifallahi2022alzheimer, ) 2022 AD Sensor+Vision Kinect depth sensor + Camera SVM 85 97.8 0.44
A-5 (jeon2023early, ) 2023 AD Sensor IMUs SVM, KNN, RF, MLP, etc 145 72.2 0.67
S-1 (xia2016novel, ) 2016 ALS Sensor Force sensitive resistor SVM 28 92.9 0.33
S-2 (felix2020effective, ) 2020 ALS Sensor Force sensitive resistor SVM, KNN. DT 29 86.2-96.6 0.33
S-3 (felix2021automatic, ) 2021 ALS Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 29 89.7 0.44
H-1 (mannini2016machine, ) 2016 HD Sensor IMUs HMM, SVM 42 88.2 0.44
H-2 (acosta2018meta, ) 2018 HD Sensor IMUs DT, RF, Logitboost, Multiboost 14 78.6-92.9 0.56
H-3 (zhang2019deep, ) 2019 HD Sensor Walkway system SVM, 3D CNN 12 82.0-86.9 0.56
H-4 (felix2019automatic, ) 2019 HD Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 36 58.3-94.4 0.44
H-5 (paula2019diagnosing, ) 2019 HD Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 36 100.0 0.67
H-6 (huang2019machine, ) 2019 HD Sensor vGRF SVM, NB, DT, RF, etc 35 85.7-97.1 0.44
H-7 (trojaniello2015assessment, ) 2020 HD Sensor
@@
M-1 (patil2019clinical, ) 2019 MS Sensor Force sensitive sensor SVM, MLP, KNN, ELM 945 89.8 0.67
M-2 (kaur2020predicting, ) 2021 MS Sensor Treadmill SVM, RF, AdaBoost, MLP, etc 40 94.3 0.44
M-3 (trentzsch2021using, ) 2021 MS Sensor Force sensitive sensor SVM, Gaussian Naive Bayes, DT, KNN, etc 60 74.5 0.44
M-4 (creagh2021interpretable, ) 2021 MS Sensor IMUs TL DCNN HAR models 97 77.6-91.1 0.89
M-5 (creagh2020smartphone, ) 2021 MS Sensor IMUs SVM, LR, RF 97 79.3-85.1 0.44
M-6 (hu2022machine, ) 2022 MS Sensor Walkway system SVM, LR, XGBoost 88 77.0-81.0 0.44
4. Overview of Existing Studies
In this section, we systematically review existing studies related to AI-assisted NDs diagnosis based on human gait. We categorize all studies by different NDs types (PD, HD, AD, ALS, MS) with a data summary for them (shown in Table 1, 2, 3, 4, and 5) and a discussion on representative studies.
Quality Evaluation of Studies. To comprehensively measure the quality of each study, we empirically grade the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples at three levels with the score 1, 2, or 3. We sum their scores and divide them by the maximum score to get a normalized quality score $S_{Q}$ ranging from 0 to 1. $S_{Q}$ is computed for each study and the detailed evaluation criterion is provided in the appendices.
4.1. Parkinson’s Disease (PD)
As shown in Table 1, 2, 3, and 4, there are 124 studies that focus on PD diagnosis based on human gait. It is worth mentioning that there are 19 studies, each of which simultaneously includes the diagnosis of PD and other NDs, as shown in Table 4 (referred to as “Combinatorial Diagnosis of Different NDs”. We separately discuss these studies in Sec. 4.2. Based on our survey, there are 77, 34, and 13 PD studies using C-ML, C-DL, and A-DL models respectively. It should be mentioned that there are a large number of PD studies (up to 124 studies), and since the studies with high-quality AI models and diagnostic results are the focus of our survey, we select representative studies from papers with quality score $S_{Q}>0.6$ for discussion.
(khorasani2014hmm, ) (Table 1, ID = P-7) utilizes Hidden Markov Models (HMM) with Gaussian Mixtures to classify gait data and differentiate Parkinson’s disease (PD) patients from healthy subjects. Using stride interval as the distinguishing feature, the HMM method achieves an accuracy rate of 90.3%. (joshi2017automatic, ) (ID = P-21) explores the use of wavelet analysis combined with SVM for the identification of PD based on spatio-temporal gait variables, achieving a classification accuracy of 90.32%. The study demonstrates the potential of using wavelet analysis as an efficient method for classifying PD and healthy subjects. In (aich2018validation, ) (ID = P-32), the authors quantify gait parameters objectively using wearable accelerometers, and compare them with a motion capture system to automatically discriminate patients with PD using machine learning models such as SVM, KNN, and DT. The proposed approach achieves an accuracy of approximately 89.1% and suggests that the wearable accelerometer-based system is suitable for assessing and monitoring PD and Freezing of gait (FOG) in real-life scenarios. (zeng2019classification, ) (ID = P-50) proposes a novel method for classifying gait patterns between patients with PD and healthy controls. This method reconstructs the phase space of vGRF and decomposes the gait dynamics using empirical mode decomposition, which are fed to neural networks (NN) for classification with accuracy of 98.8%. The study of (som2020unsupervised, ) (ID = P-60) explores the use of single-sensor accelerometer data from healthy subjects to extract features relevant to multi-sensor accelerometer gait data for Parkinson’s disease classification. They use a pre-trained convolutional autoencoder as the source model and a simple multi-layer perceptron as the target model to classify PD and achieve 68.92% accuracy. In (ghaderyan2021inter, ) (ID = P-81), a time-varying singular value decomposition method is designed to extract the most useful and informative part of the vGRF signal data, and then sparse non-negative least squares (NNLS) approach is leveraged to classify PD and HCs, achieving 97.2% accuracy. The study of (reyes2019lstm, ) (ID = P-43) leverages convolutional LSTM to identify PD patients from Kinect-based skeleton data and achieves up to 83% accuracy. It also suggests that the data preprocessing ($e.g.$, cropped data) can help boost the classification performance. (el2020deep, ) (ID = P-52) devises a deep neural network (DNN) based model to identify Parkinsonian gait based on vGRF data and achieves 98.7% accuracy. This is also the first study to predict the Unified Parkinson’s Disease Rating Scale (UPDRS) severity ($i.e.$, Parkinsion severity rate) of a subject and achieve 85.3% accuracy. (carvajal2022classification, ) (ID = P-86) combines CNN and gated recurrent units (GRUs) to perform classification between PD and health subjects from IMU signals, achieving 83.7-92.7% accuracy. The results suggest that the combination of temporal and spectral information is more effective in classifying gait patterns of PD patients. (alharthi2020gait, ) (ID = P-75) uses deep CNN to classify gait deterioration in PD patients based on spatio-temporal vGRF signals and achieves 95.5% accuracy in average. The study suggests that the heel strike and body balance are the most indicative gait elements in PD classification. (salimi2023type, ) (ID = P-85) proposes an interval type-2 fuzzy neural network with novel quasi-Levenberg-Marquardt (qLM) learning approach to identify PD patients from the gait cycle based vGRF signals. Based on the 10 clinical features extracted from 14 gait features, the model achieves 97.6% accuracy with a smaller size of parameters and a good interpretability ($i.e.$, rules based on fuzzy logic). (zhang2023wm, ) (ID = P-121) proposes a novel model named WM-STGCN that leverages a Weighted adjacency matrix and Multi-scale temporal convolution in a spatio-temporal Graph Convolution Network to classify PD gaits from normal gaits and achieves up to 87.1% accuracy on 2D skeleton data. This method outperforms many existing methods including LSTM, KNN, DT, AdaBoost, and ST-GCN models (yan2018spatial, ). (dong2023static, ) (ID = P-123) devises a static-dynamic temporal networks that consists of one-dimensional, two-dimensional convolutional networks, and attention mechanisms to classify PD patients based on foot force (vGRF time series) and achieves 96.7% accuracy.
4.2. Combinatorial Diagnosis of Different NDs
@@
As shown in Table 4, there are 19 studies that simultaneously focus on diagnosis of PD, ALS, and HD based on human gait, and 11, 6, 2 studies use C-ML, C-DL, A-DL models, respectively. We provide a brief content summary and discussion for representative studies below.
@@
(dutta2013hybrid, ) (ID = C-2) uses a cross-correlation method to extract features with recurrent neural networks to classify HC and people with neurodegenerative diseases (PD, HD, ALS), achieving highly-competitive performance with 90.6% accuracy. The study of (xia2015classification, ) (ID = C-3) exploits 9 statistical features (mean, minimum, etc) of gait rhythm signals ($i.e.$, gait cycle parameters) to perform neurodegenerative disease (ND) diagnosis using different classic machine learning models (SVM, RF, MLP, KNN), and achieves 96.6-100% accuracy. Najafabadian $etal.$ (najafabadian2018neurodegenerative, ) (ID = C-5) present a nonlinear gait signal analysis approach to classify NDDs by extracting chaotic and fractal features from gait data and reach a classification accuracy of 92.34%. (athisakthi2017statistical, ) (ID = C-6) introduces a technique called Statistical Energy Values and Peak Analysis (SEP) for detecting NDDs from signals obtained through force-sensitive resistors. The accuracy, sensitivity, and specificity values are best achieved using the Random Forest Classifier reported at 96.9%, 96.8%, and 96.7% respectively. (yan2020classification, ) (ID = C-8) introduces a framework called Topological Motion Analysis (TMA) which embeds gait fluctuation time series into a phase space. In detail, persistent homology is used to extract topological signatures of barcodes, and a Random Forest classifier is employed for analysis. This study demonstrates the effectiveness of TMA in classifying diseases such as ALS, HD, and PD compared to HC, achieving high accuracy of 79.31%, 91.67%, and 87.10% respectively. (lin2020evaluation, ) (ID = C-11) proposes a CNN model to differentiate gait of NDs patients and HC based on vGRF signal data (10-second signal) and achieves 95.95-100% accuracy. (beyrami2020robust, ) (ID = C-12) introduces a non-invasive, cost-effective diagnostic tool for analysing gait signals obtained from vGRF sensors. The best classification results are obtained using SVM model and the achieved average accuracy rates are 100% for ALS, 99.78% for PD, and 99.90% for HD. A hybrid model that combines multiple sensor data is proposed in (zhao2021multimodal, ) (ID = C-15) to learn gait differences between ALS, HD, PD, and healthy individuals. The model includes a spatial feature extractor (SFE) and a correlative memory neural network (CorrMNN) for capturing representative features and temporal information, which outperforms many state-of-the-art techniques with 99.7-100.0% accuracy. (tobar2022petri, ) (ID = C-17) devise Petri net models of human gait with Rs and classification trees (CTs) to classify different neurodegenerative diseases (PD, HD, ALS) from Petri net features of vGRFs, and achieves 91.7-95.1% accuracy.
4.3. Huntington’s Disease (HD)
@@
(mannini2016machine, ) (Table 5, ID = H-1) proposes a method to classify gait between Huntinton’s disease (HD) and health elderly based on IMU data using HMM derived features and SVM classifier, and achieves overall accuracy 90.5% by combining HMM-based features, time and frequency domain features. In (acosta2018meta, ) (ID = H-2), 11 gait features are extracted from accelerometer data to classify HD patients using different machine learning models (RF, DT, Logiboost, Multibooost) and achieves up to 92.86% accuracy. Two attribute selection algorithms are proposed to choose the most representative gait patterns of subjects. (zhang2019deep, ) uses pressure data collected during formation of individual footsteps to classify HD by 3D CNN. The experiments showed that using the basic deep learning backbone VGG16 and similar modules can achieve classification accuracy of 89%. Different machine learning models (KNN, SVM, NB, DT, LDA) are utilized in (felix2019automatic, ) to identify HD from foot force based gait dynamics parameters and achieves up to 100% accuracy with SVM and DT when considering the right foot stance interval. In (huang2019machine, ), five machine learning models (SVM, DT, NV, RF, LR) are utilized to recognize HD from gait dynamics features and achieves 85.71-97.14% accuracy. (trojaniello2015assessment, ) uses the IMU data and SVM model to classify HD gait and healthy gait of elderly, achieving 90.5% accuracy.
4.4. Multiple Sclerosis (MS)
@@
In (patil2019clinical, ) (Table 5, ID = M-1), the authors uses Extreme Learning Machine (ELM), SVM, MLP, and KNN to classify multiple gait disorders including MS, stroke, and cerebral palsy from gait force data, achieving a accuracy result of 89.8%. The study of (creagh2021interpretable, ) (ID = M-4) devises a CNN-based transfer learning framework to identify MS patients from wearable smartphone sensor data and achieves up to 91.1% accuracy. The interpretations of results also suggest that cadence-based measures, gait speed, and ambulation-related signal perturbations are distinct features for classification. (hu2022machine, ) (ID = M-6) leverages gait parameters extracted from raw data of a walkway system to identify HD patients using different classic machine learning models (SVM, LR, XGBoost) and achieves up to 81% accuracy. The performance is shown to be further improved to 88% when augmenting standard parameters with other custom parameters and normalized subject characteristics.
4.5. Alzheimer’s Disease (AD)

### v2

@@ 4.2.3 A-DL Models for PD Diagnosis
C-19 [220] 2023 PD, ALS, HD Sensor Force sensitive resistor LSTM, GRU, ResNet, FCN, etc 64 75.4-83.3 0.56
Table 7: Existing studies of AI-assisted diagnosis of AD (A-1 to A-5), ALS (S-1 to S-3), HD (H-1 to H-7) or MS (M-1 to M-6) based on gait. We summarize input data, main AI model(s), experimental sample size, accuracy (range), and quality score of each study. Quality scores that evaluate the novelty of AI models, comprehensiveness of method comparison, and sufficiency of experimental samples are computed (see Sec. 4.1).
@@
A-1 [221] 2016 AD Sensor IMUs SVM, MLP, RBNs, DBSs 72 91.0-96.6 0.56
A-2 [71] 2020 AD Sensor+Vision Camera+EEG SVM, RF, CNN, ST-GCN, etc 87 91.1-98.6 0.89
A-3 [222] 2020 AD Sensor Pressure sentsitive mat SVM 78 78.0 0.44
A-4 [107] 2022 AD Sensor+Vision Kinect depth sensor + Camera SVM 85 97.8 0.44
A-5 [223] 2023 AD Sensor IMUs SVM, KNN, RF, MLP, etc 145 72.2 0.67
S-1 [104] 2016 ALS Sensor Force sensitive resistor SVM 28 92.9 0.33
S-2 [224] 2020 ALS Sensor Force sensitive resistor SVM, KNN. DT 29 86.2-96.6 0.33
S-3 [225] 2021 ALS Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 29 89.7 0.44
H-1 [48] 2016 HD Sensor IMUs HMM, SVM 42 88.2 0.44
H-2 [226] 2018 HD Sensor IMUs DT, RF, Logitboost, Multiboost 14 78.6-92.9 0.56
H-3 [227] 2019 HD Sensor Walkway system SVM, 3D CNN 12 82.0-86.9 0.56
H-4 [228] 2019 HD Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 36 58.3-94.4 0.44
H-5 [229] 2019 HD Sensor Force sensitive resistor SVM, KNN, NB, LDA, etc 36 100.0 0.67
H-6 [230] 2019 HD Sensor vGRF SVM, NB, DT, RF, etc 35 85.7-97.1 0.44
H-7 [231] 2020 HD Sensor
@@
M-1 [105] 2019 MS Sensor Force sensitive sensor SVM, MLP, KNN, ELM 945 89.8 0.67
M-2 [59] 2021 MS Sensor Treadmill SVM, RF, AdaBoost, MLP, etc 40 94.3 0.44
M-3 [232] 2021 MS Sensor Force sensitive sensor SVM, Gaussian Naive Bayes, DT, KNN, etc 60 74.5 0.44
M-4 [233] 2021 MS Sensor IMUs TL DCNN HAR models 97 77.6-91.1 0.89
M-5 [106] 2021 MS Sensor IMUs SVM, LR, RF 97 79.3-85.1 0.44
M-6 [58] 2022 MS Sensor Walkway system SVM, LR, XGBoost 88 77.0-81.0 0.44
4.3 Combinatorial Diagnosis of Different NDs
@@
As shown in Table 6 and 1, there are 19 studies that simultaneously focus on diagnosis of PD, ALS, and HD based on human gait, and 11, 6, 2 studies use C-ML, C-DL, A-DL models, respectively. We provide a brief content summary and discussion for representative studies below.
@@
[101] (Table 6, ID = C-1) uses force sensitive sensors to collect stride intervals and footfall contact times. Features are later derived from these time series data and filtered using a genetic algorithm for diagnosing purpose. The best accuracy results for HC $vs.$ HD, HC $vs.$ PD, HC $vs.$ ALS are achieved using the features selected and SVM models with the corresponding rate of 90.3%, 89.3%, and 96.8% respectively while the accuracy for the diagnosis of three NDDs altogether is 90.6%. [206] (ID = C-2) uses a cross-correlation method to extract features with recurrent neural networks to classify HC and people with neurodegenerative diseases (PD, HD, ALS), achieving highly-competitive performance with 90.6% accuracy. The study of [102] (ID = C-3) exploits 9 statistical features (mean, minimum, etc) of gait rhythm signals ($i.e.$, gait cycle parameters) to perform neurodegenerative disease (ND) diagnosis using different classic machine learning models (SVM, RF, MLP, KNN), and achieves 96.6-100% accuracy. [207] (ID = C-4) proposes a novel analytical methodology for distinguishing healthy individuals and NDs patients based on gait patterns. The method employs intrinsic mode functions (IMFs), and extracts discriminative features from the area measure and relative change in centroid position of the polygon formed by the Convex Hull of these analytic IMFs. Najafabadian $etal.$ [208] (ID = C-5) present a nonlinear gait signal analysis approach to classify NDDs by extracting chaotic and fractal features from gait data and reach a classification accuracy of 92.34%. [209] (ID = C-6) introduces a technique called Statistical Energy Values and Peak Analysis (SEP) for detecting NDDs from signals obtained through force-sensitive resistors. The accuracy, sensitivity, and specificity values are best achieved using the Random Forest Classifier reported at 96.9%, 96.8%, and 96.7% respectively. In [210] (ID = C-7), the authors transform left foot stride and stride interval time series data into a string using symbolic aggregate approximation (SAX). The converted string is then utilized to find a string prototype for each disease using the proposed algorithm called sgUPFCMed. Finally, they implements a fuzzy KNN model to find the best match for a test data sample to classify different diseases and HCs. [211] (ID = C-8) introduces a framework called Topological Motion Analysis (TMA) which embeds gait fluctuation time series into a phase space. In detail, persistent homology is used to extract topological signatures of barcodes, and a Random Forest classifier is employed for analysis. This study demonstrates the effectiveness of TMA in classifying diseases such as ALS, HD, and PD compared to HC, achieving high accuracy of 79.31%, 91.67%, and 87.10% respectively. [212] (ID = C-9) utilizes two feature extraction approaches, $i.e.$, Recurrence Quantification Analysis (RQA) and Fast Walsh-Hadamard Transform (FWHT), to extract gait features and then feed them into different machine learning models such as RF, SVM, and KNN to classify NDs and HC. [213] (ID = C-10) leverages differential transformation of gait force (GF) signals and multiscale sample entropy (MSE) to classify different neurodegenerative diseases using KNN and SVM. This study also shows that the accuracy can be improved by applying the method SMOTE to balance the amount of data in each class. [214] (ID = C-11) proposes a CNN model to differentiate gait of NDs patients and HC based on vGRF signal data (10-second signal) and achieves 95.95-100% accuracy. [103] (ID = C-12) introduces a non-invasive, cost-effective diagnostic tool for analysing gait signals obtained from vGRF sensors. The best classification results are obtained using SVM model and the achieved average accuracy rates are 100% for ALS, 99.78% for PD, and 99.90% for HD. [215] (ID = C-13) explores the usefulness of state space vector, which is combined with KNN, to learn features from gait data for NDD detection. Besides, it also involves statistical features’ extraction from gait signals. It also shows the state space approach’s potential for analyzing other physiological signals. [216] (ID = C-14) builds a preprocessing process, a feature transformation process (PCA), and a classification process (CNN and SVM) to leverage five-minute time-frequency spectrogram of vGRF to classify neurodegenerative diseases. A hybrid model that combines multiple sensor data is proposed in [77] (ID = C-15) to learn gait differences between ALS, HD, PD, and healthy individuals. The model includes a spatial feature extractor (SFE) and a correlative memory neural network (CorrMNN) for capturing representative features and temporal information, which outperforms many state-of-the-art techniques with 99.7-100.0% accuracy. [217] (ID = C-16) proposes to use wavelet coherence to extract features from gait force signal and to represent those extracted features by using spectrogram image. Those features are then processed by CNN to classify HCs and NDs. [218] (ID = C-17) devise Petri net models of human gait with Rs and classification trees (CTs) to classify different neurodegenerative diseases (PD, HD, ALS) from Petri net features of vGRFs, and achieves 91.7-95.1% accuracy. Spatio-temporal gait data are used in [219] (ID = C-18) to measure the energy cost and power spectral density in gait pathologies of different NDs, while fuzzy c-means techniques, learning algorithm for multivariate data analysis (LAMDA), and neural networks are exploited to classify between the neurodegenerative diseases and the control group. In [220] (ID = C-19), the authors compare various deep learning models including LSTM, GRU, InceptionTime, ResNet, FCN, TST, and PatchTST in their ability to differentiate healthy gaits from PD, ALS, and HD gaits.
4.4 Huntington’s Disease (HD)
@@
[48] (Table 7, ID = H-1) proposes a method to classify gait between Huntinton’s disease (HD) and health elderly based on IMU data using HMM derived features and SVM classifier, and achieves overall accuracy 90.5% by combining HMM-based features, time and frequency domain features. In [226] (ID = H-2), 11 gait features are extracted from accelerometer data to classify HD patients using different machine learning models (RF, DT, Logiboost, Multibooost) and achieves up to 92.86% accuracy. Two attribute selection algorithms are proposed to choose the most representative gait patterns of subjects. [227] uses pressure data collected during formation of individual footsteps to classify HD by 3D CNN. The experiments showed that using the basic deep learning backbone VGG16 and similar modules can achieve classification accuracy of 89%. Different machine learning models (KNN, SVM, NB, DT, LDA) are utilized in [228] to identify HD from foot force based gait dynamics parameters and achieves up to 100% accuracy with SVM and DT when considering the right foot stance interval. [229] leverages classic machine learning models (SVM, DT, KNN, NB, LDA) to identify HD from the extracted gait features such as right stance interval. In [230], five machine learning models (SVM, DT, NV, RF, LR) are utilized to recognize HD from gait dynamics features and achieves 85.71-97.14% accuracy. [231] uses the IMU data and SVM model to classify HD gait and healthy gait of elderly, achieving 90.5% accuracy.
4.5 Multiple Sclerosis (MS)
@@
In [105] (Table 7, ID = M-1), the authors uses Extreme Learning Machine (ELM), SVM, MLP, and KNN to classify multiple gait disorders including MS, stroke, and cerebral palsy from gait force data, achieving a accuracy result of 89.8%. [59] (ID = M-2) proposes two normalization methods to normalized the subject’s derived gait characteristics, which are fed into ML models such as SVM, RF, Adaboost, and MLP to classified MS and HC. [232] (ID = M-3) investigates the most efficient gait parameters that could aid in the detection of subtle gait changes in individuals with Multiple Sclerosis. Aiming at this goal, it compares three diagnostic gait systems and employs six different machine learning algorithms. It contributes to the field by providing insights into the use of technology and machine learning in the management of MS. The study of [233] (ID = M-4) devises a CNN-based transfer learning framework to identify MS patients from wearable smartphone sensor data and achieves up to 91.1% accuracy. The interpretations of results also suggest that cadence-based measures, gait speed, and ambulation-related signal perturbations are distinct features for classification. In [106] (ID = M-5), the authors leverage smart-phone and smart-watch to collect gait data and compute related features. Then, the features are feed as input to the ML models (SVM, LR, RF) to classify HC and MS. [58] (ID = M-6) leverages gait parameters extracted from raw data of a walkway system to identify HD patients using different classic machine learning models (SVM, LR, XGBoost) and achieves up to 81% accuracy. The performance is shown to be further improved to 88% when augmenting standard parameters with other custom parameters and normalized subject characteristics.
4.6 Alzheimer’s Disease (AD)